# ImFace++: A Sophisticated Nonlinear 3D Morphable Face Model with Implicit Neural Representations

### v2

@@ 3 Method
Our goal is to learn a neural network to parameterize $f$, making it satisfy the genuine facial shape priors. As shown in Fig. 2, the proposed network for ImFace++ is constructed by a series of Mini-Nets blocks that are progressively learned. In the first stage, ImFace++ employs the Expression and Identity Mini-Nets blocks to learn separate deformation fields associated with expression- and identity-variation, respectively. Meanwhile, a Template Mini-Nets block is used to learn a signed distance field of a template face shape. Such explicitly disentangled architecture ensures that the intricate expression deformations and inter-individual differences can be accurately modeled (Sec. 3.1). The second stage complements the former by incorporating a Detail Mini-Nets block to learn a refinement displacement field (RDF) within the template space, which faithfully encodes facial geometry details (see Sec. 3.2). The network parameters and facial embeddings are jointly learned. After convergence, a face scan be disentangled by deforming it from its observed state to the template.
All the aforementioned fields are implemented by a shared Mini-Nets architecture, where the entire facial deformation or geometry is decomposed into an array of semantically meaningful parts and encoded by a set of local field functions so that intricate details can be sufficiently captured. A lightweight module conditioned on the query point position, i.e, Fusion Network, is stacked at the end of the Mini-Nets block to adaptively blend the local fields. As such, an elaborate Neural Blend-Field is achieved (see Sec. 3.3). Additionally, an improved embedding learning strategy is designed, allowing more diverse details of expression (refer to Sec. 3.4). Detailed explanations of the model designs are provided as follows.
@@ 3.1 Disentangled Deformation Fields
The proposed method initially focuses on non-linear facial deformation modeling and the establishment of correspondences among different faces. This is accomplished through three Mini-Nets blocks as previously outlined. As Fig. 2 (a) shows, we collectively refer to this process as Disentangled Deformation Fields modeling and the learned SDF can reproduce a base observation space. The core components serve distinct purposes, with slight structural adaptations as follows:
@@ 3.1.1 Expression Mini-Nets (ExpNet)
The facial deformations incurred by expressions are represented by ExpNet $\mathcal{E}$, which learns an observation-to-canonical warping for every face scan:
@@
where ${\mathbf{p}}_{b}$ denotes a query point in the base observation space, and $l\in{\mathbb{R}}^{k\times 3}$ denotes $k$ 3D landmarks on the observed face generated by a Landmark-Net $\eta:({\mathbf{z}}_{exp},{\mathbf{z}}_{id})\mapsto l$, which is introduced to localize the query point ${\mathbf{p}}_{b}$ in the Neural Blend-Field. This point is deformed by $\mathcal{E}$ to ${\mathbf{p}}^{\prime}$ in the person-specific canonical space, which represents faces with a neutral expression.
@@ 3.1.2 Identity Mini-Nets (IDNet)
To model shape morphs among individuals, IDNet $\mathcal{I}$ further warps the canonical space to a template shape space shared by all faces:
@@
where $l^{\prime}\in{\mathbb{R}}^{k\times 3}$ denotes $k$ landmarks on the canonical face generated by another Landmark-Net conditioned only on the identity embedding, namely $\eta^{\prime}:{\mathbf{z}}_{id}\mapsto l^{\prime}$, where ${\mathbf{p}}^{\prime\prime}$ is the deformed point in the template space. To cope with the possible non-existent correspondences produced during preprocessing, $\mathcal{I}$ additionally predicts a residual term $\delta\in{\mathbb{R}}$ to correct the SDF value $s_{0}$ predicted by the subsequent TempNet, which is similar to [23].
@@ 3.1.3 Template Mini-Nets (TempNet)
$\hat{f}({\mathbf{p}}_{b},{\mathbf{z}}_{exp},{\mathbf{z}}_{id})=\mathcal{T}(\mathcal{I}(\mathcal{E}({\mathbf{p}}_{b})))+\mathcal{I}_{\delta}(\mathcal{E}({\mathbf{p}_{b}})),$ (5)
For brevity, the latent embeddings and predicted landmarks are omitted. By doing so, we can automatically learn correspondence across different individuals by mapping the query point ${\mathbf{p}}_{b}$ to the template point ${\mathbf{p}^{\prime\prime}}$.
Figure 3: Neural Blend-Field. (a) The Mini-Nets block is a shared architecture, which decomposes an entire facial feature into semantically meaningful parts and encodes them by a set of local field functions. It is tailed by a Fusion Network for more comprehensive representations. (b) The Landmark-Net is introduced to softly partition the entire facial surface. (c) The Fusion Network is a lightweight module conditioned on the query point position, which adaptively blends the local field functions, resulting in an elaborate Neural Blend-Field.
3.2 Refinement Displacement Field within Template Space
Facial modeling assumes a pivotal role in enhancing the overall visual fidelity and realism of facial representations. The intricate nuances, such as fine lines, wrinkles, and subtle expressions, collectively contribute to the creation of a truly lifelike and convincing depiction of a face. In order to construct an advanced 3DMM, we have introduced a specialized refinement module designed to master high-frequency facial details. Drawing inspiration from the concept of Implicit Displacement Fields (IDF) proposed in [82], where high-frequency details are generated by offsetting a smooth base surface along its normal directions based on the displacement distances predicted from IDF, we have developed a Detail Mini-Nets block (DetailNet), denoted as $\mathcal{D}$, to learn a refinement displacement field (RDF), as depicted in Figure 2 (b). To account for the dynamic nature of human faces, instead of representing facial geometry details using a conventional IDF, we encode them within the individual-independent template space that exhibits relative stability, so that subtle changes can be well captured. This strategy is reminiscent of the concept of storing displacement information in a topology-consistent UV texture map.
To achieve this, we leverage the correspondences established during the Disentangled Deformation Fields modeling phase and RDF is learned in the later stage. Formally, its modeling process can be expressed as:
@@ 3.2 Refinement Displacement Field within Template Space
where ${\mathbf{d}}$ denotes the displacement distance along the normal of an observed point ${\mathbf{p}}$, which tends to be overlooked in the first learning stage. It is worth noting that ${\mathbf{p}}^{\prime\prime}$ is the input of this function, ensuring that the subtle geometry details are learned in the template space. Such learning phase decomposition effectively isolates subtle displacements from identity and expression-related deformations, and the facial details can be readily applied to other facial shapes that share the same template space. Taking the detail displacement into consideration, the coordinates ${\mathbf{p}}_{b}$ in the aforementioned coarse base observation space represented by $\hat{f}$ can be corrected as follows:
@@
where $\chi$ denotes the attenuation function [82] and $\frac{\nabla\hat{f}({\mathbf{p}})}{||\nabla\hat{f}({\mathbf{p}})||}$ denotes the normalized normal of ${\mathbf{p}}$. Finally, for a face scan, its key components (i.e., expression, and identity) can be accurately disentangled by performing the following warping process:
@@
The proposed ImFace++ learns face morphs by disentangled deformation fields and RDF in a fine-grained and meaningful manner, ensuring that more diverse and sophisticated facial deformations, correspondence, and details can be accurately learned.
@@ 3.3 Neural Blend-Field
$v={\bm{\psi}}({\mathbf{x}})=\sum_{n=1}^{k}{{\it w}_{n}({\mathbf{x}}){\bm{\psi}}_{n}({\mathbf{x}}-l_{n})},\vspace{-2mm}$ (9)
where $l_{n}$ is a parameter that describes the $n$-th local region, ${\it w}_{n}({\mathbf{x}})$ is the $n$-th blend weight, ${\bm{\psi}}_{n}({\mathbf{x}}-l_{n})$ is the corresponding local field, and $k$ is the number of local regions. In this way, the blending is performed on a series of local fields, rather than calculating a weighted average of the output values $v$ of some fixed positions, leading to more powerful representation capability in handling complicated local features.
@@ 3.3.1 Deformation Formulation
We formulate the deformation with a SE(3) field $({\bm{\omega}},{\mathbf{v}})\in{\mathbb{R}}^{6}$, where ${\bm{\omega}}\in so(3)$ is a rotate vector representing the screw axis and the angle of rotation. The deformed coordinates ${\mathbf{x}}^{\prime}$ can be calculated by ${e^{\bm{\omega}}}{\mathbf{x}}+{\mathbf{t}}$, where the rotation matrix ${e^{\bm{\omega}}}$ (exponential map form of Rodrigues’ formula) is written as:
@@ 3.3.2 Hyper Nets
To obtain a more compact and expressive latent space, we have introduced a meta-learning approach[45]. A Hyper Net $\phi_{n}$ is implemented by an MLP and it predicts the instance-specific parameters for ExpNet $\mathcal{E}$, IDNet $\mathcal{I}$ and DetailNet $\mathcal{D}$. It takes a latent code ${\mathbf{z}}$ as input and generates the parameters for the neurons in a Mini-Net ${\bm{\psi}}_{n}$ so that the learned facial representations possess a higher variety.
@@
Figure 4: (a) SDF is able to represent closed shapes. (b) UDF is capable of representing an open surface but the gradient is discontinuous at the boundary, making it hard to be fitted by neural networks. (c) The proposed method generates pseudo watertight faces and restricts implicit functions on them, enabling implicit neural networks to learn geometry representations on 3D faces.
@@ 3.5 Loss Functions
where ${\bar{s}}$ and ${\mathbf{\bar{n}}}$ denote the ground-truth SDF values and the field gradients, respectively. $i$ denotes the index of sampled faces, and $\lambda$ indicates trade-off parameter. ${\Omega_{i}}$ is the sampling space of the scan $i$.
@@
where ${\cal L}_{eik}^{i}$ serves the purpose of enabling the network to satisfy the Eikonal constraint[26] in both the observation and canonical spaces concurrently. This constraint plays a significant role in maintaining a sensible correspondence throughout facial deformations at all stages of the network.
@@
In the initial training phase, our focus lies on training the disentangled deformation fields, as elaborated in Section 3.1. This phase involves learning the coarse base SDF and establishing correspondences between various facial shapes. The training loss is formulated as follows:
@@
Subsequently, we freeze the sub-networks $\mathcal{E}$, $\mathcal{I}$, and $\mathcal{T}$ and proceed to train $\mathcal{D}$ to learn the refinement displacement field, by substituting $\hat{f}$ with $f$ in the following training loss:
@@ 4 Experiments
We conduct thorough subjective and objective evaluations of ImFace++, along with ablation studies to validate the effectiveness of the specifically designed modules. Finally, we demonstrate various applications of ImFace++.
@@ 4.1 Dataset
FaceScape[41] is utilized for evaluation. It is a large-scale high quality 3D face dataset consisting of 938 individuals with 20 types of expressions. The data from 847 individuals are publicly accessible and we mainly use them for experiments. More specifically, 16,434 face scans from 831 individuals with nearly 20 expressions are sampled as the training set, and another set of 319 face scans from the remaining 16 individuals with nearly 20 expressions are used as our testing set. It is worth noting that, before conducting any data preprocessing, we make sure that the face scans include high-frequency details by making use of the displacement map provided by FaceScape.
@@
Figure 5: Illustration of the preprocessing pipeline. The original face data from FaceScape [41] undergoes initial normalization and cropping. Hidden surfaces are subsequently removed. The Delaunay Triangulation Algorithm [50] is then employed for achieving watertight results. Lastly, we sample points and estimate the SDF and normals for training ImFace++.
@@
Data Preprocessing. Given that neural networks are particularly adept at fitting functions that are differentiable across the entire domain, recent studies on implicit functions generally require watertight input data. While functions like UDF do not strictly necessitate watertightness, they exhibit non-differentiability when crossing a surface and may struggle to effectively handle fine details, as illustrated in Figure 4. To address this, we have developed an efficient preprocessing pipeline that transforms the input data into pseudo-watertight faces. This preprocessing step enables the establishment of a general SDF on these faces. By doing so, we ensure that the geometry and correspondence are learned with the same level of precision and fidelity as that of watertight objects.
@@
As illustrated in Figure 5, the preprocessing consists of a series of steps: i) We rigidly align the facial scans to a frontal orientation using facial landmarks. Each mesh is then normalized to a standard unit of 10 cm. ii) The coordinate origin is defined at a point 4 cm behind the nose tip. A sphere with a radius of 10 cm is set as the sampling area, denoted as $S$. Mesh triangles outside of this sphere are cropped. iii) We apply the Ray-Triangle Intersection Algorithm [51] to remove hidden surfaces, such as the nasal and oral cavities. iv) Delaunay Triangulation[50] is performed on the x-y coordinates of the mesh to create an oriented and pseudo-watertight structure. v) Since FaceScape provides 3D face scans with uniform topology, we sample 13,008 points as landmarks within the defined sampling area $S$ of each scan to train ImFace++.
@@
With the pseudo-watertight facial structures generated, we compute the SDF values through a distance transform. The sign of each sample is determined by the angle between its distance vector to the nearest surface and the positive z-axis direction. Values behind the facial surface are defined as negative. To ensure comprehensive coverage, we uniformly sample 250,000 points on each facial surface: 125,000 points near the facial surface with random displacements sampled from a normal distribution $\mathcal{N}(0,1cm)$, and 125,000 points within the sphere. For each of these points, we calculate their signed distances and gradient vectors. In the end, the sampled data is organized into triplets, denoted as ${({\bf p},{\bf\bar{n}},\bar{s})}$, representing the query point, gradient vectors, and signed distance value. These triplets serve as the input data for training ImFace++.

### v3

@@ 3 Method
Our goal is to learn a neural network to parameterize $f$, making it satisfy the genuine facial shape priors. As shown in Fig. 2, the proposed network for ImFace++ is constructed by a series of Mini-Nets blocks that are progressively learned. In the first stage, ImFace++ employs the Expression and Identity Mini-Nets blocks to learn separate deformation fields associated with expression- and identity-variations, respectively. Meanwhile, a Template Mini-Nets block is used to learn a signed distance field of a template face shape. Such explicitly disentangled architecture ensures that the intricate expression deformations and inter-individual differences can be accurately modeled (Sec. 3.1). The second stage complements the former by incorporating a Detail Mini-Nets block to learn a refinement displacement field, i.e. RDF, within the template space, which faithfully encodes facial geometry details (see Sec. 3.2). The network parameters and facial embeddings are jointly learned. After convergence, a face scan can be disentangled by deforming it from its detailed state to the template.
All the aforementioned fields are implemented by a shared Mini-Nets architecture, where the entire facial deformation or geometry is decomposed into an array of semantically meaningful parts and encoded by a set of local field functions so that intricate details can be sufficiently captured. A lightweight module conditioned on the query point position, i.e, Fusion Network, is stacked at the end of the Mini-Nets block to adaptively blend the local fields. As such, an elaborate Neural Blend-Field is achieved (see Sec. 3.3). Additionally, an improved embedding learning strategy is designed, allowing more diverse details of expressions (refer to Sec. 3.4). Detailed explanations of the model designs are provided as follows.
@@ 3.1 Disentangled Deformation Fields
The proposed method initially focuses on non-linear facial deformation modeling and establishment of point-to-point correspondences among different faces. This is accomplished through three Mini-Nets blocks as previously outlined. As Fig. 2 (a) shows, we collectively refer to this process as Disentangled Deformation Fields modeling and the learned SDF can finally reproduce an expression space (Fig. 9 illustrates the differences between these spaces), which encompasses facial shapes with different expressions and identities. The core components serve distinct purposes, with slight structural adaptations as follows:
@@ 3.1.1 Expression Mini-Nets (ExpNet)
The facial deformations incurred by expressions are represented by ExpNet $\mathcal{E}$, which learns a warping from the expression space to the identity space for each face scan:
@@
where ${\mathbf{p}}_{b}$ denotes a query point in the expression space, and $l\in{\mathbb{R}}^{k\times 3}$ denotes $k$ 3D landmarks on faces in the expression space generated by a Landmark-Net $\eta:({\mathbf{z}}_{exp},{\mathbf{z}}_{id})\mapsto l$, which is introduced to localize the query point ${\mathbf{p}}_{b}$ in the Neural Blend-Field. This point is deformed by $\mathcal{E}$ to ${\mathbf{p}}^{\prime}$ in the person-specific identity space, which represents faces with a neutral expression.
@@ 3.1.2 Identity Mini-Nets (IDNet)
To model shape morphs among individuals, IDNet $\mathcal{I}$ further warps the identity space to a template shape space shared by all faces:
@@
where $l^{\prime}\in{\mathbb{R}}^{k\times 3}$ denotes $k$ landmarks on a face with a neutral expression generated by another Landmark-Net conditioned only on the identity embedding, namely $\eta^{\prime}:{\mathbf{z}}_{id}\mapsto l^{\prime}$, where ${\mathbf{p}}^{\prime\prime}$ is the deformed point in the template space. To cope with the possible non-existent correspondences produced during preprocessing, $\mathcal{I}$ additionally predicts a residual term $\delta\in{\mathbb{R}}$ to correct the SDF value $s_{0}$ predicted by the subsequent TempNet, which is similar to [23].
@@ 3.1.3 Template Mini-Nets (TempNet)
$\hat{f}({\mathbf{p}}_{b},{\mathbf{z}}_{exp},{\mathbf{z}}_{id})=\mathcal{T}(\mathcal{I}(\mathcal{E}({\mathbf{p}}_{b})))+\mathcal{I}_{\delta}(\mathcal{E}({\mathbf{p}_{b}})).$ (5)
For brevity, the latent embeddings and predicted landmarks are omitted. $\mathcal{I}_{\delta}(\mathcal{E}({\mathbf{p}_{b}}))$ denotes the residual term $\delta$ in Eq. (3). By doing so, we can automatically learn correspondence across different individuals by mapping the query point ${\mathbf{p}}_{b}$ to the template point ${\mathbf{p}^{\prime\prime}}$.
Figure. 3: Neural Blend-Field. (a) The Mini-Nets block is a shared architecture, which decomposes an entire facial feature into semantically meaningful parts and encodes them by a set of local field functions. It is tailed by a Fusion Network for more comprehensive representations. (b) The Landmark-Net is introduced to softly partition the entire facial surface. (c) The Fusion Network is a lightweight module conditioned on the query point position, which adaptively blends the local field functions, resulting in an elaborate Neural Blend-Field.
3.2 Refinement Displacement Field
Facial modeling assumes a pivotal role in enhancing the overall visual fidelity and realism of facial representations. The intricate nuances, such as fine lines, wrinkles, and subtle expressions, collectively contribute to the creation of a truly lifelike and convincing depiction of a face. To construct an advanced 3DMM, we introduce a specialized refinement module designed to master high-frequency facial details. Drawing inspiration from the concept of Implicit Displacement Fields (IDF) proposed in [82], where high-frequency details are generated by offsetting a smooth base surface along its normal directions based on the displacement distances predicted from IDF, we develop a Detail Mini-Nets block (DetailNet), denoted as $\mathcal{D}$, to learn an RDF, as depicted in Fig. 2 (b). To account for the dynamic nature of human faces, instead of representing facial geometry details using a conventional IDF, we encode them in the individual-independent template space that exhibits a relative stability so that subtle changes can be well captured. This strategy is reminiscent of the concept of storing displacement information in a topology-consistent UV texture map.
To achieve this, we leverage the correspondences established during disentangled deformation field modeling and RDF is learned in the later stage. Formally, its modeling process can be expressed as:
@@ 3.2 Refinement Displacement Field
where ${\mathbf{d}}$ denotes the displacement distance along the normal of a point ${\mathbf{p}}$, which tends to be overlooked in the first learning stage. It is worth noting that ${\mathbf{p}}^{\prime\prime}$ is the input of this function, ensuring that the subtle displacements are learned in the template space. Such learning phase decomposition effectively isolates subtle displacements from identity and expression-related deformations, and the facial details can be readily applied to other facial shapes that share the same template space. Taking the detail displacement into consideration, the coordinates of ${\mathbf{p}}_{b}$ in the aforementioned expression space represented by $\hat{f}$ can be corrected as follows:
@@
where $\chi$ denotes the attenuation function [82] and $\mathbf{p}^{\prime\prime}=\mathcal{I}(\mathcal{E}(\mathbf{p}_{b}))$. $\frac{\nabla\hat{f}({\mathbf{p}})}{||\nabla\hat{f}({\mathbf{p}})||}$ denotes the normalized normal of ${\mathbf{p}}$ calculated in the expression space. This approach accounts for the influence of expressions when modeling facial details and ensures that geometries not captured by the disentangled deformation fields are fully learned. Finally, for a face scan, its key components (i.e., expression and identity) can be accurately disentangled by performing the following warping process:
@@
Notably, while Eq. (8) ideally requires $\mathbf{p}_{b}$ as input, obtaining exact matching pairs ($\mathbf{p}_{b}$, $\mathbf{p}$) in the 3D space is difficult as the facial shapes are typically provided in explicit form like meshes. For practical implementation, we approximate $\mathbf{p}^{\prime\prime}=\mathcal{I}(\mathcal{E}(\mathbf{p}))$, considering the small displacement distance. The proposed ImFace++ learns face morphs by disentangled deformation fields and RDF in a fine-grained and meaningful manner, ensuring that more diverse and sophisticated facial deformations, correspondence, and details can be accurately captured.
@@ 3.3 Neural Blend-Field
$v={\bm{\psi}}({\mathbf{x}})=\sum_{n=1}^{k}{{\it w}_{n}({\mathbf{x}}){\bm{\psi}}_{n}({\mathbf{x}}-l_{n})},$ (9)
where $l_{n}$ is a parameter that describes the $n$-th local region, ${\it w}_{n}({\mathbf{x}})$ is the $n$-th blend weight, ${\bm{\psi}}_{n}({\mathbf{x}}-l_{n})$ is the corresponding local field, and $k$ is the number of local regions. In this way, the blending is performed on a series of local fields, rather than calculating a weighted average of the output values $v$ of some fixed positions, leading to a more powerful representation capability in handling complicated local features.
@@ 3.3.1 Deformation Formulation
We formulate the deformation with an SE(3) field $({\bm{\omega}},{\mathbf{v}})\in{\mathbb{R}}^{6}$, where ${\bm{\omega}}\in so(3)$ is a rotate vector representing the screw axis and the angle of rotation. The deformed coordinates ${\mathbf{x}}^{\prime}$ can be calculated by ${e^{\bm{\omega}}}{\mathbf{x}}+{\mathbf{t}}$, where the rotation matrix ${e^{\bm{\omega}}}$ (exponential map form of Rodrigues’ formula) is written as:
@@ 3.3.2 Hyper Nets
To obtain a more compact and expressive latent space, we introduce a meta-learning approach [45]. A Hyper Net $\phi_{n}$ is implemented by an MLP and it predicts the instance-specific parameters for ExpNet $\mathcal{E}$, IDNet $\mathcal{I}$ and DetailNet $\mathcal{D}$. It takes a latent code ${\mathbf{z}}$ as input and generates the parameters for the neurons in a Mini-Net ${\bm{\psi}}_{n}$ so that the learned facial representations possess a higher variety.
@@
Figure. 4: (a) SDF is able to represent closed shapes. (b) UDF is capable of representing an open surface but the gradient is discontinuous at the boundary, making it hard to be fitted by neural networks. (c) The proposed method generates pseudo watertight faces and restricts implicit functions on them, enabling implicit neural networks to learn geometry representations on 3D faces.
@@ 3.5 Loss Functions
where ${\bar{s}}$ and ${\mathbf{\bar{n}}}$ denote the ground-truth SDF values and the field gradients, respectively. $i$ denotes the index of sampled faces and $\lambda$ indicates the trade-off parameter. ${\Omega_{i}}$ is the sampling space of the scan $i$.
@@
where ${\cal L}_{eik}^{i}$ serves the purpose of enabling the network to satisfy the Eikonal constraint[26] in both the expression and identity spaces concurrently. This constraint ensures that both the spaces adhere to the SDF properties, where the gradient norm is uniformly unit, thereby preserving the physical integrity of the deformation field.
@@
In the initial training phase, our focus lies on training the disentangled deformation fields, as elaborated in Sec. 3.1. This phase involves learning the coarse base SDF and establishing correspondences between various facial shapes. The training loss is formulated as follows:
@@
Subsequently, we freeze the sub-networks $\mathcal{E}$, $\mathcal{I}$, and $\mathcal{T}$ and proceed to train $\mathcal{D}$ to learn the RDF, by substituting $\hat{f}$ with $f$ in the following training loss:
@@ 4 Experiments
We conduct thorough subjective and objective evaluation of ImFace++, along with ablation studies to validate the effectiveness of the specifically designed modules. Finally, we demonstrate various applications of ImFace++.
@@ 4.1 Dataset
FaceScape [41] is used for evaluation. It is a large-scale high quality 3D face dataset consisting of 938 individuals with 20 types of expressions. The data from 847 individuals are publicly accessible and we mainly adopt them for experiments. More specifically, 16,434 face scans from 831 individuals are sampled as the training set, and 319 face scans from the remaining 16 individuals are utilized as the testing set, where all the types of expressions are almost included. It is worth noting that, before conducting any data preprocessing, we make sure that the face scans include high-frequency details by making use of the displacement map provided by FaceScape. Additionally, 100 face scans from the LYHM [104] dataset are employed for cross-dataset evaluation.
@@
Figure. 5: Illustration of the preprocessing pipeline. The original face data from FaceScape [41] undergo initial normalization and cropping. Hidden surfaces are subsequently removed. The Delaunay Triangulation algorithm [50] is then employed to achieve watertight results. Lastly, we sample points and estimate the SDF and normals for training ImFace++.
@@
Data Preprocessing. Given that neural networks are particularly adept at fitting functions that are differentiable across the entire domain, recent studies on implicit functions generally require watertight input data. While functions like UDF do not strictly necessitate watertightness, they exhibit a non-differentiability when crossing a surface and may struggle to handle fine details, as illustrated in Fig. 4. To address this, we develop an efficient preprocessing pipeline that transforms the input data into pseudo-watertight faces. This preprocessing step enables the establishment of a general SDF on these faces. By doing so, we ensure that the geometry and correspondence are learned with the same level of precision and fidelity as that of watertight objects.
@@
As shown in Fig. 5, the preprocessing pipeline consists of a series of steps: i) We rigidly align the facial scans to a frontal orientation using facial landmarks. Each mesh is then normalized to a standard unit of 10 cm. ii) The coordinate origin is defined at a point 4 cm behind the nose tip. A sphere with a radius of 10 cm is set as the sampling area. Mesh triangles outside of this sphere are cropped. iii) We apply the Ray-Triangle Intersection algorithm [51] to remove hidden surfaces, such as the nasal and oral cavities. iv) Delaunay Triangulation[50] is performed on the x-y coordinates of the mesh to create an oriented and pseudo-watertight structure. v) Since FaceScape provides 3D face scans with a uniform topology, we sample 13,008 points as landmarks within the defined sampling area of each scan to train ImFace++.
@@
With the pseudo-watertight facial structures generated, we compute the SDF values through a distance transform. The sign of each sample is determined by the angle between its distance vector to the nearest surface and the positive z-axis direction. The values behind the facial surface are defined as negative. To ensure comprehensive coverage, we uniformly sample 250,000 points on each facial surface: 125,000 points near the facial surface with random displacements sampled from a normal distribution $\mathcal{N}(0,1cm)$, and 125,000 points within the sphere. For each of these points, we calculate their signed distances and gradient vectors. In the end, the sampled data are organized into triplets, including the query point, gradient vectors, and signed distance value. These triplets serve as the input data for training ImFace++.